OpenStudy (anonymous):

Kim is studying the sale of a particular brand of cereals from the year 2000 to 2010. She writes the following function to model the sale of the cereal S(t), in million dollars, after t years: S(t) = t2 + 7t + 69 Part A: What does the y-intercept of the graph of the function represent? (4 points) Part B: What is the reasonable domain of the graph of the function? (3 points) Part C: What is the average rate of change of the sale of the cereal from the first year to the fourth year? Show your work. (3 points) I JUST NEED PART C!!!!! PLEASEEEEEE JUST PART C IF ANYBODY CAN HELP ME

OpenStudy (anonymous):

@campbell_st

OpenStudy (anonymous):

average rate of change will be the slope from the points of year 1 and year 4, so, lets calculate what the sale is for year 1, what do you get?

OpenStudy (anonymous):

@DemolisionWolf 2000

OpenStudy (anonymous):

@DemolisionWolf is it 2000

OpenStudy (anonymous):

S(t) = t2 + 7t + 69 let t = 1 = 1^2 + 7(1) + 69 = 77 so, we have a point of (1, 77) now u try with year 4 so, t=4

OpenStudy (anonymous):

@DemolisionWolf so for year 1 it is 4 years. Can you help me with the year 2??

OpenStudy (anonymous):

so the total sales for year 1, was $77 to find the total sales for year 4, we need to plug in 4 every where we see 't' in the equation: t^2 + 7t + 69 and solve

OpenStudy (anonymous):

@DemolisionWolf So when i do 4*4 and get my answer that will be the answer for year 4??

OpenStudy (anonymous):

@DemolisionWolf Can you stil help me??

OpenStudy (anonymous):

umm i'm not sure i understand why you want to do 4*4? to get the answer for year 4, just do t=4 t^2 + 7t + 69 and solve

OpenStudy (anonymous):

@DemolisionWolf How do I do that!!!!!

OpenStudy (anonymous):

@DemolisionWolf I got 76.

OpenStudy (anonymous):

t^2 + 7t + 69 4^2 + 7(4) + 69 16 + 28 + 69